OpenStudy (anonymous):

The second differential, d^2y/dx^2 of a curve is -3x+1 (1) Determine the expression of the gradient function, if the gradient at x=0 is 4. (11) the expression for the equation of the curve if it passes through (1,0)

OpenStudy (anonymous):

How do you make vertical bars in latex?

OpenStudy (anonymous):

\[\frac{d^2y}{dx^2} = -3x + 1 \] The gradient of a curve is \(\frac{dy}{dx}\) so integrate the second differential for the gradient

OpenStudy (anonymous):

Vertical bars..?

OpenStudy (anonymous):

and for the question \(x=0\) is given; then easy to solve for \(C\)

OpenStudy (anonymous):

like... \[\frac{dy}{dx}|_{x=t}\]except the bar is the same size as the derivative.

OpenStudy (anonymous):

i mean at x = 0 at 4*

OpenStudy (anonymous):

Sorry i dont know herp_derp

OpenStudy (amistre64):

you have to incase it in left right syntax \left. stuff \right|

OpenStudy (amistre64):

\[\left. \frac{top}{bottom} \right|_{down}^{up}\]